OpenStudy (anonymous):

Which of the following statements is false? A. 3x^2 + 4x^2 = 7x^4 B. 3x^2 • 4x^2 = 12x^4 C. (3x^3)^2 = 9x^6 D. –3^2 = –9

OpenStudy (anonymous):

Is D the correct answer?

OpenStudy (goldphenoix):

Yes.

OpenStudy (goldphenoix):

Why? Well because 3*2 is equal to what?

OpenStudy (goldphenoix):

OH WAIT. NO.

OpenStudy (goldphenoix):

I thought it said 3*2. >_>

OpenStudy (goldphenoix):

-3^2 is another way of saying -3 * -3. 3 *3 = 9 A negative multiply a negative is a positive. So it's a positive 9 or just 9.

OpenStudy (anonymous):

So which statement is false?

OpenStudy (goldphenoix):

Guess and check.

OpenStudy (goldphenoix):

Let's try C. C. (3x^3)^2 = 9x^6 What does C. (3x^3)^2 become?

OpenStudy (anonymous):

9x^5?

OpenStudy (goldphenoix):

Well let see. Another way of writing that is: C. (3x^3)(3x^3) 3 * 3 = 9 x^3 * x^3= x^6

OpenStudy (goldphenoix):

So that is true. Though that's not the correct answer. Let's try B: B. 3x^2 • 4x^2 = 12x^4 What does 3x^2 • 4x^2 give you?

OpenStudy (anonymous):

D is right for sure

OpenStudy (anonymous):

D is also right \[(3x^3)^2 = 9x^6\] is true

OpenStudy (anonymous):

the wrong one is the first one \[ 3x^2 + 4x^2 = 7x^4\]

OpenStudy (anonymous):

for example, if \(x=10\) then \(300+400=700\) not \(300+400=70000\)

OpenStudy (goldphenoix):

So isn't the answer either A or D?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

only A is wrong

OpenStudy (anonymous):

I am so confused right now!

OpenStudy (anonymous):

why is A wrong?

OpenStudy (anonymous):

@GoldPhenoix you made a mistake here -3^2 is another way of saying -3 * -3.

OpenStudy (anonymous):

in fact \(-3^2=-3\times 3=-9\)

OpenStudy (anonymous):

A is wrong because when you add, you do not add exponents

OpenStudy (goldphenoix):

Really?! I never knew! o_o

OpenStudy (anonymous):

3 apples and 4 apples is 7 apples 3 x squared and 4 x squared is 7 x squared

OpenStudy (anonymous):

that is what i was trying to explain when i said " put \(x=10\)" if \(x=10\) then \(3x^2=300\) and \(4x^2=400\) and when you add you get \[700=7\times 10^2\] not \(7\times 10^4\)

OpenStudy (anonymous):

in other words, \[3x^2+4x^2=7x^2\]

OpenStudy (anonymous):

@GoldPhenoix you did so know that suppose you were asked to compute \[50-3^2\] what would you get?

OpenStudy (goldphenoix):

50-9=41

OpenStudy (anonymous):

right, so you knew that \(-3^2=-9\) or else you would have said \[50-3^2=50+9=59\] which is silly

OpenStudy (goldphenoix):

Oh, I see.

OpenStudy (anonymous):

so A is wrong...

OpenStudy (goldphenoix):

Therefore, A is the correct answer in your question.

OpenStudy (anonymous):

yeah... thanks!:)